OpenStudy (anonymous):

it says solve the salloing equations write the variables above their valures to find the name of a famous volcano a+18= -32??

OpenStudy (anonymous):

its -50!

OpenStudy (anonymous):

steps?

OpenStudy (anonymous):

-32 - 18 = a

OpenStudy (anonymous):

is that setp by step?

OpenStudy (anonymous):

madhatter u need help in simple terms

OpenStudy (anonymous):

Yeah its really simple

OpenStudy (anonymous):

see we have a + 18 = -32 which is constant term and which one is variable

OpenStudy (anonymous):

-32-18=A but i need the steps ?

OpenStudy (anonymous):

that is the steps.... you switch a and -32 and then you switch the plus sign to the opposite which is a minus.

OpenStudy (anonymous):

those are the steps**

OpenStudy (anonymous):

a + 18 = -32 subtract 18 from both sides a+ 18 -18 = -32 -18 a + 0 = -50 a = -50

OpenStudy (anonymous):

mad how old are u?

OpenStudy (anonymous):

lol yeah could you help me find what z is please now??

OpenStudy (anonymous):

yeah sure go ahead post it

OpenStudy (anonymous):

sheg xcan u help me with 4 more?

OpenStudy (anonymous):

yeah post it somewhere else <-----------------------------